OpenStudy (anonymous):

Tammy’s Diner donated 6% of their weekend profit to charity. If the weekend profit of the Diner was $5,268, how much money was donated to charity?

OpenStudy (anonymous):

LOL! Still here to help!

OpenStudy (anonymous):

$5,268 x .06 = $314.08

OpenStudy (anonymous):

You have to basically do 5268 multiplied by 6%, and there is your answer!

OpenStudy (anonymous):

100% weekend profit: 5,268 6% to donate: x (6 * 5,268) / 100 = x (31608) / 100 = x 316.08= x

OpenStudy (anonymous):

Got it? good luck yet again!

OpenStudy (anonymous):

I apologize for the typo in my earlier response. It is $316.08

OpenStudy (anonymous):

:D Thx for the medal!

OpenStudy (anonymous):

Michelle is salesperson at a retail store. She earns 9% commission on her total sales. If Michelle’s total sales were $9,860, how much money did she earn?